A cruise ship can sail 32 mph in calm water. Sailing with the Gulf Stream, the ship can sail 228 mi in the same amount of time as it takes to sail 156 mi against the Gulf Stream. Find the rate of the Gulf Stream.
Answer by mananth(16949)

A cruise ship can sail 32 mph in calm water. Sailing with the Gulf Stream, the ship can sail 228 mi in the same amount of time as it takes to sail 156 mi against the Gulf Stream. Find the rate of the Gulf Stream.
...
rate of current be xmph
..
rate with current = 32+x : d= 228 miles
rate against current = 32-x: d=156 miles
..
times are same, so equate the time
..
228/(32+x)= 156/(32-x)
228x(32-x)=156(32+x)
7296-228x= 4992+156x
reorganise
228x+156x=7296-4992
72x= 2304
/72
x= 2304/72
x= 32 mph the rate of current
...
m.ananth@hotmail.ca